OpenStudy (anonymous):

The distance from the top of the archery is 1.5 m if you release at a height of 1.39m and the arrow hits 18m away what is the quadratic equation?

OpenStudy (anonymous):

ax^2+bx+c=0

OpenStudy (anonymous):

\[-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

got it? @hayhay5270

OpenStudy (anonymous):

No not really lol I tried but I cannot figure out the a b or c to solve the equation

OpenStudy (anonymous):

@Brostep0s

OpenStudy (anonymous):

are these your points? (0, 1.39) (18,40) and (45,0).

OpenStudy (anonymous):

I just have the facts and my measurement that's it I have no clue

OpenStudy (anonymous):

well, those should be your points, if my calculations are correct.

OpenStudy (anonymous):

Okay yeah so I'm just stuck on that part of finding the maximum height and the parabolic path

OpenStudy (anonymous):

use this for your equation: 0 = a(45)^2 +b(45) +1.39

OpenStudy (anonymous):

should come out to 1.41 if you do everything correctly.

OpenStudy (anonymous):

How would I solve that? I'm not to good at algebra

OpenStudy (anonymous):

a=45 b=45 c=1.39 \[-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

Okay that's whatis got so 1.41 is my max height ?

OpenStudy (anonymous):

yeppers

OpenStudy (anonymous):

Okay thanks!

OpenStudy (anonymous):

cx

OpenStudy (anonymous):

;)